The preferred stock of McKinney lines , inc, pays an annual dividend of $6.45 and sells for$75.70 a share what is the rate of return on this sicurity?
Answer #1

Stock price=(Annual dividend)/(Required rate of return)
Substituting the values, we get;
75.70=(6.45)/(Required rate of return)
(Required rate of return)=(6.45)/75.70
=>(Required rate of return)=(6.45)/75.70=0.085204756 or 8.52% (Rounded to two decimal places)